# A general framework of Riemannian adaptive optimization methods with a convergence analysis

### v1

@@ 1 Introduction
Bonnabel (2013) proposed Riemannian stochastic gradient descent (RSGD), the most basic Riemannian stochastic optimization algorithm. In particular, Riemannian stochastic variance reduction algorithms, such as Riemannian stochastic variance-reduced gradient (RSVRG) (Zhang et al., 2016), Riemannian stochastic recursive gradient (RSRG) (Kasai et al., 2018), and Riemannian stochastic path-integrated differential estimator (R-SPIDER) (Zhang et al., 2018; Zhou et al., 2019), are based on variance reduction methods in Euclidean space. There are several prior studies on Riemannian adaptive optimization methods for specific Riemannian manifolds. In particular, Kasai et al. (2019) proposed a Riemannian adaptive stochastic gradient algorithm on matrix manifolds (RASA). RASA is an adaptive optimization method on matrix manifolds (e.g., the Stiefel manifold or the Grassmann manifold), with a convergence analysis under the upper-Hessian bounded and retraction $L$-smooth assumptions (see (Kasai et al., 2019, Section 4) for details). However, RASA is not a direct extension of the adaptive optimization methods commonly used in deep learning, and it works only for diminishing step sizes. On the cartesian product of Riemannian manifolds, RAMSGrad (Bécigneul & Ganea, 2019) and modified RAMSGrad (Sakai & Iiduka, 2021), direct extensions of AMSGrad, have been proposed as methods that work on Cartesian products of Riemannian manifolds. In particular, Roy et al. (2018) proposed cRAMSProp and applied it to several Riemannian stochastic optimizations. However, they did not provide a convergence analysis of cRAMSProp. More recently, Riemannian stochastic optimization methods, Sharpness-aware minimization on Riemannian manifolds (Riemannian SAM) (Yun & Yang, 2024) and Riemannian natural gradient descent (RNGD) (Hu et al., 2024), were proposed.
@@ 1.1 Contributions
Motivated by the above discussion, we propose a framework of adaptive optimization methods on Riemannian submanifolds of Euclidean space (Algorithm 1) that is based on the framework (Reddi et al., 2018, Algorithm 1) proposed by Reddi, Kale and Kumar for Euclidean space. Our framework incorporates the mini-batch strategy that is often used in deep learning. Important examples of Riemannian submanifolds of the Euclidean space include the unit sphere and the Stiefel manifold. Moreover, within this framework, we propose AMSGrad on embedded submanifolds of Euclidean space (Algorithm 2) as a direct extension of AMSGrad. In addition, we give convergence analyses (Theorem 3.7) valid for both a constant step size (Theorem 3.8) and diminishing step size (Theorem 3.9). Our analyses not only ensure that the proposed method converges to the optimal solution, but also reveal the relationship between the convergence rate and mini-batch size. Moreover, we numerically compare the performances of several methods based on Algorithm 1, including Algorithm 2, with the existing methods. In the numerical experiments, we applied the algorithms to principal component analysis (PCA) (Kasai et al., 2018; Roy et al., 2018) and the low-rank matrix completion (LRMC) problem (Boumal & Absil, 2015; Kasai et al., 2019; Hu et al., 2024), which can be considered to be Riemannian optimization problems.
Our first contribution is to propose a general framework of Riemannian adaptive optimization methods (Algorithm 1) and AMSGrad on embedded submanifolds of Euclidean space (Algorithm 2). In particular, the proposed method incorporates the mini-batch strategy. Our second contribution is to give convergence analyses of Algorithms 1 and 2. In particular, we emphasize that the proposed method can use both constant and diminishing step sizes (Theorems 3.8 and 3.9), in contrast to RASA (Kasai et al., 2019), which only uses a diminishing step size. The third contribution is to compare the proposed methods with RSGD and RASA in numerical experiments.
@@ 2 Mathematical Preliminaries
Let $\mathbb{R}^{d}$ be a $d$-dimensional Euclidean space with inner product $\left\langle{x,y}\right\rangle_{2}:=x^{\top}y$, which induces the norm $\left\lVert{\cdot}\right\rVert_{2}$. Let $\mathbb{R}_{++}$ be the set of positive real numbers, i.e., ++:=$\mathbb{R}_{++}:=\{x\in\mathbb{R}\mid x>0\}$. $I_{d}$ denotes a $d\times d$ identity matrix. For square matrices $X,Y\in\mathbb{R}^{d\times d}$, we write $X\prec Y$ (resp. $X\preceq Y$) if $Y-X$ is a positive-definite (resp. positive-semidefinite) matrix. For two matrices $X$ and $Y$ of the same dimension, $X\odot Y$ denotes the Hadamard product, i.e., element-wise product. Let $\max(X,Y)$ be the element-wise maximum. Let $\mathcal{S}^{d}$ (resp. $\mathcal{S}^{d}_{+}$, $\mathcal{S}^{d}_{++}$) be the set of $d\times d$ symmetric (resp. symmetric positive-semidefinite, symmetric positive-definite) matrices, i.e., $S^{d}:=\{X\in\mathbb{R}^{d\times d}\mid X^{\top}=X\}$, $\mathcal{S}^{d}_{+}:=\{X\in\mathbb{R}^{d\times d}\mid X\succeq O\}$ and $\mathcal{S}^{d}_{++}:=\{X\in\mathbb{R}^{d\times d}\mid X\succ O\}$. Let $\mathcal{D}^{d}$ be the set of $d\times d$ diagonal matrices. Let $\mathcal{O}_{d}$ be the orthogonal group, i.e., $\mathcal{O}_{d}:=\{X\in\mathbb{R}^{d\times d}\mid X^{\top}X=I_{d}\}$.
Let $M$ be an embedded submanifold of $\mathbb{R}^{d}$. Moreover, let $T_{x}M$ be the tangent space at a point $x\in M$ and $TM$ be the tangent bundle of $M$. Let $0_{x}$ be the zero element of $T_{x}M$. The inner product $\left\langle{\cdot,\cdot}\right\rangle_{2}$ of a Euclidean space $\mathbb{R}^{d}$ induces a Riemannian metric $\left\langle{\cdot,\cdot}\right\rangle_{x}$ of $M$ at $x\in M$ according to $\left\langle{\xi,\eta}\right\rangle_{x}=\left\langle{\xi,\eta}\right\rangle_{2}=\xi^{\top}\eta$ for $\xi,\eta\in T_{x}M\subset T_{x}\mathbb{R}^{d}\cong\mathbb{R}^{d}$. The norm of $\eta\in T_{x}M$ is defined as $\left\lVert{\eta}\right\rVert_{x}=\sqrt{\eta^{\top}\eta}=\left\lVert{\eta}\right\rVert_{2}$. Let $P_{x}:T_{x}\mathbb{R}^{d}\cong\mathbb{R}^{d}\to T_{x}M$ be the orthogonal projection onto $T_{x}M$ (see Absil et al. (2008)). For a smooth map $F:M\to N$ between two manifolds $M$ and $N$, $\mathrm{D}F(x):T_{x}M\to T_{F(x)}N$ denotes the derivative of $F$ at $x\in M$. The Riemannian gradient $\operatorname*{grad}f(x)$ of a smooth function $f:M\to\mathbb{R}$ at $x\in M$ is defined as a unique tangent vector at $x$ satisfying $\left\langle{\operatorname*{grad}f(x),\eta}\right\rangle_{x}=\mathrm{D}f(x)[\eta]$ for any $\eta\in T_{x}M$.
@@ 2.1 Examples
The unite sphere $\mathbb{S}^{d-1}:=\{x\in\mathbb{R}^{d}\mid\left\lVert{x}\right\rVert_{2}=1\}$ is an embedded manifold of $\mathbb{R}^{d}$. The tangent space $T_{x}\mathbb{S}^{n-1}$ at $x\in\mathbb{S}^{d-1}$ is given by $T_{x}\mathbb{S}^{n-1}=\{\eta\in\mathbb{R}^{d}\mid\eta^{\top}x=0\}$. The induced Riemannian metric on $\mathbb{S}^{d-1}$ is given by $\left\langle{\xi,\eta}\right\rangle_{x}=\left\langle{\xi,\eta}\right\rangle_{2}:=\xi^{\top}\eta$ for $\xi,\eta\in T_{x}\mathbb{S}^{d-1}$. The orthogonal projection $P_{x}:\mathbb{R}^{d}\to T_{x}\mathbb{S}^{d-1}$ onto the tangent space $T_{x}\mathbb{S}^{n-1}$ is given by $P_{x}(\eta)=(I_{d}-xx^{\top})\eta$ for $x\in\mathbb{S}^{d-1}$ and $\eta\in T_{x}\mathbb{S}^{d-1}$.
@@
The induced Riemannian metric on $\operatorname*{St}(p,n)$ is given by $\left\langle{\xi,\eta}\right\rangle_{X}=\operatorname*{tr}(\xi^{\top}\eta)$ for $\xi,\eta\in T_{X}\operatorname*{St}(p,n)$. The orthogonal projection onto the tangent space $T_{X}\operatorname*{St}(p,n)$ is given by $P_{X}(\eta)=\eta-X\operatorname*{sym}(X^{\top}\eta)$ for $X\in\operatorname*{St}(p,n)$, $\eta\in T_{X}\operatorname*{St}(p,n)$, where $\operatorname*{sym}(A):=(A+A^{\top})/2$. The Stiefel manifold $\operatorname*{St}(p,n)$ reduces to the orthogonal groups when $n=p$, i.e. $\operatorname*{St}(p,p)=\mathcal{O}_{p}$.

### v2

@@ 1 Introduction
Bonnabel (2013) proposed Riemannian stochastic gradient descent (RSGD), the most basic Riemannian stochastic optimization algorithm. In particular, Riemannian stochastic variance reduction algorithms, such as Riemannian stochastic variance-reduced gradient (RSVRG) (Zhang et al., 2016), Riemannian stochastic recursive gradient (RSRG) (Kasai et al., 2018), and Riemannian stochastic path-integrated differential estimator (R-SPIDER) (Zhang et al., 2018; Zhou et al., 2019), are based on variance reduction methods in Euclidean space. There are several prior studies on Riemannian adaptive optimization methods for specific Riemannian manifolds. In particular, Kasai et al. (2019) proposed a Riemannian adaptive stochastic gradient algorithm on matrix manifolds (RASA). RASA is an adaptive optimization method on matrix manifolds (e.g., the Stiefel manifold or the Grassmann manifold), and gave a convergence analysis under the upper-Hessian bounded and retraction $L$-smooth assumptions (see (Kasai et al., 2019, Section 4) for details). However, RASA is not a direct extension of the adaptive optimization methods commonly used in deep learning, and it works only for diminishing step sizes. RAMSGrad (Bécigneul & Ganea, 2019) and modified RAMSGrad (Sakai & Iiduka, 2021), direct extensions of AMSGrad, have been proposed as methods that work on Cartesian products of Riemannian manifolds. In particular, Roy et al. (2018) proposed cRAMSProp and applied it to several Riemannian stochastic optimizations. However, they did not provide a convergence analysis of it. More recently, Riemannian stochastic optimization methods, sharpness-aware minimization on Riemannian manifolds (Riemannian SAM) (Yun & Yang, 2024) and Riemannian natural gradient descent (RNGD) (Hu et al., 2024), were proposed.
1.1 Motivations
Kasai et al. (2019) presented useful results indicating RASA with a diminishing step size converges to a stationary point of a continuously differentiable function defined on a matrix manifold. Meanwhile, other results have indicated that using constant step sizes makes the algorithm useful for training deep neural networks (Zaheer et al., 2018) and that the setting of the mini-batch size depends on the performance of the algorithms used to train deep neural networks (Smith et al., 2018; Sato & Iiduka, 2023). The main motivation of this paper is thus to gain an understanding of the theoretical performance relationship between Riemannian adaptive optimization methods using constant step sizes and ones using the mini-batch size. We are also interested in understanding the theoretical performance relationship between the Riemannian adaptive optimization methods using diminishing step sizes and those using the mini-batch size.
@@ 1.2 Contributions
Our first contribution is to propose a framework of adaptive optimization methods on Riemannian submanifolds of Euclidean space (Algorithm 1) that is based on the framework (Reddi et al., 2018, Algorithm 1) proposed by Reddi, Kale and Kumar for Euclidean space. A key challenge in designing adaptive gradient methods over Riemannian manifolds is the absence of a coordinate system like that in Euclidean space. In this work, we specifically focus on manifolds that are embedded in Euclidean space, a common setting in many applications. We address the issue by constructing algorithms that rely on projections onto the tangent space at each point on the manifold. Our framework incorporates the mini-batch strategy that is often used in deep learning. Important examples of Riemannian submanifolds of the Euclidean space include the unit sphere and the Stiefel manifold. Note that (Kasai et al., 2019) focuses exclusively on the RASA algorithm, which is defined only on matrix manifolds. In contrast, our framework can be defined on any manifold embedded in Euclidean space. Moreover, within this framework, we propose AMSGrad on embedded submanifolds of Euclidean space (Algorithm 2) as a direct extension of AMSGrad.
Our second contribution is to give convergence analyses of the proposed algorithm (Algorithm 1) by using a mini-batch size $b$ valid for both a constant step size (Theorem 3.5) and diminishing step size (Theorem 3.6). In particular, Theorem 3.5 indicates that, under some conditions, the sequence $(x_{k})_{k=1}^{\infty}$ generated by the proposed algorithm with a constant step size $\alpha$ satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{K}+\frac{1}{b}\right).$
We should note that Theorem 3.5 is an extension of the result in Zaheer et al. (2018, Corollary 2). Theorem 3.6 indicates that, under some conditions, the sequence $(x_{k})_{k=1}^{\infty}$ generated by the proposed algorithm with a diminishing step size $\alpha_{k}$ satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\left(1+\frac{1}{b}\right)\frac{\log K}{\sqrt{K}}\right).$
Since the stochastic gradient (see (1)) using large mini-batch sizes is approximately the full gradient of $f$; intuitively, using large mini-batch sizes decreases the number of steps needed to find stationary points of $f$. Theorems 3.5 and 3.6 indicate that, the larger the mini-batch size $b$ is, the smaller the upper bound of $\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}[\|\mathrm{grad}f(x_{k})\|_{2}^{2}]$ becomes and the fewer the required steps become. Hence, Theorems 3.5 and 3.6 would match our intuition.
Theorem 3.5 indicates that, although the proposed algorithm with a constant step size $\alpha$ and constant mini-batch size $b$ does not always converge, we can expect that converges with an increasing mini-batch size $b_{k}$ such that $b_{k}\leq b_{k+1}$. In practice, it has been shown that increasing the mini-batch size (Byrd et al., 2012; Balles et al., 2017; De et al., 2017; Smith et al., 2018; Goyal et al., 2018) is useful for training deep neural networks with mini-batch optimizers. Motivated by Theorems 3.5 and 3.6, we give convergence analyses of the proposed algorithm by using an increasing mini-batch size $b_{k}$ valid for both a constant step size (Theorem 3.7) and diminishing step size (Theorem 3.8). Theorem 3.7 indicates that, under some conditions, the sequence $(x_{k})_{k=1}^{\infty}$ generated by the proposed algorithm with a constant step size $\alpha$ and an increasing mini-batch size $b_{k}$ satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{K}\right).$
That is, with increasing mini-batch sizes, it can find a stationary point of $f$, in contrast to Theorem 3.5. Moreover, Theorem 3.8 indicates that, under some conditions, the sequence $(x_{k})_{k=1}^{\infty}$ generated by the proposed algorithm with a diminishing step size $\alpha_{k}$ and an increasing mini-batch size $b_{k}$ satisfies
$\min_{k=1,\ldots,K}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]\leq\frac{1}{\sum_{k=1}^{K}\alpha_{k}}\sum_{k=1}^{K}\alpha_{k}\mathbb{E}\left[\left\lVert{\operatorname*{grad}f(x_{k})}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(\frac{1}{\sqrt{K}}\right).$
That is, with increasing mini-batch sizes, it has a better convergence rate than with constant mini-batch sizes (Theorem 3.6).
The third contribution is to numerically compare the performances of several methods based on Algorithm 1, including Algorithm 2, with the existing methods. In the numerical experiments, we applied the algorithms to principal component analysis (PCA) (Kasai et al., 2018; Roy et al., 2018) and the low-rank matrix completion (LRMC) problem (Boumal & Absil, 2015; Kasai et al., 2019; Hu et al., 2024), which can be considered to be Riemannian optimization problems. Based on our experiments, we can summarize the following key findings. For the PCA problem, RAMSGrad with a constant step size minimizes the objective function in fewer or comparable iterations compared with RSGD, RASA-L, RASA-R and RASA-LR, regardless of the dataset. For the LRMC problem, RAdam with a diminishing step size minimizes the objective function in fewer or about the same number of iterations compared with RSGD, RASA-L and RASA-R, irrespective of the dataset. As reported in Sakai & Iiduka (2021), RAMSGrad demonstrates robust performance regardless of the initial step size. For the PCA and LRMC problems (where LRMC adopts a diminishing step size), a grid search was conducted, allowing the selection of an initial step size that yielded even better results. This likely contributed to the strong performance of RAMSGrad across datasets. However, for the LRMC problem, RAMSGrad may underperform relative to RASA, depending on the dataset. It is worth noting that the objective function in the LRMC problem is complex, and assumptions such as retraction $L$-smoothness may not hold in this case. This could explain the less favorable results observed for RAMSGrad in this particular setting. Moreover, to support our theoretical analyses, we show that increasing the batch size decreases the number of steps needed to a certain index. In particular, Table 1 (resp. Table 2) supports the result shown in Theorem 3.5, demonstrating that the proposed algorithms with a constant (resp. diminishing) step size require fewer iterations to reduce the gradient norm below the threshold as the batch size increases. Tables 1 and 2 confirm that larger batch sizes lead to faster convergence in terms of the number of iterations, aligning with our theoretical analysis. Similar tables (Tables 1116) for the remaining datasets and experiments can be found in Appendix E. Note that the choice of batch size plays a critical role in balancing the computational cost per iteration and the total number of iterations required for convergence. This is related to the concept of the critical batch size, as discussed in Sakai & Iiduka (2024), which minimizes a stochastic first-order oracle (SFO). However, estimating this optimal batch size is challenging due to the dependency on unknown parameters such as the gradient Lipschitz constant and the variance of the stochastic gradient. To address this challenge, an effective approach may involve dynamically increasing the batch size at regular intervals during optimization, such as doubling or tripling it after a fixed number of iterations. This method has theoretical guarantees for convergence (see Theorems 3.7 and 3.8) and can adaptively approach an efficient balance between computational cost and convergence speed.
@@ 2 Mathematical preliminaries
Let $\mathbb{R}^{d}$ be a $d$-dimensional Euclidean space with inner product $\left\langle{x,y}\right\rangle_{2}:=x^{\top}y$, which induces the norm $\left\lVert{\cdot}\right\rVert_{2}$. Let $\mathbb{R}_{++}$ be the set of positive real numbers, i.e., ++:=$\mathbb{R}_{++}:=\{x\in\mathbb{R}\mid x>0\}$. $I_{d}$ denotes a $d\times d$ identity matrix. For square matrices $X,Y\in\mathbb{R}^{d\times d}$, we write $X\prec Y$ (resp. $X\preceq Y$) if $Y-X$ is a positive-definite (resp. positive-semidefinite) matrix. For two matrices $X$ and $Y$ of the same dimension, $X\odot Y$ denotes the Hadamard product, i.e., element-wise product. Let $\max(X,Y)$ be the element-wise maximum. Let $\mathcal{S}^{d}$ (resp. $\mathcal{S}^{d}_{+}$, $\mathcal{S}^{d}_{++}$) be the set of $d\times d$ symmetric (resp. symmetric positive-semidefinite, symmetric positive-definite) matrices, i.e., $S^{d}:=\{X\in\mathbb{R}^{d\times d}\mid X^{\top}=X\}$, $\mathcal{S}^{d}_{+}:=\{X\in\mathbb{R}^{d\times d}\mid X\succeq O\}$ and $\mathcal{S}^{d}_{++}:=\{X\in\mathbb{R}^{d\times d}\mid X\succ O\}$. Let $\mathcal{D}^{d}$ be the set of $d\times d$ diagonal matrices. Let $\mathcal{O}_{d}$ be the orthogonal group, i.e., $\mathcal{O}_{d}:=\{X\in\mathbb{R}^{d\times d}\mid X^{\top}X=I_{d}\}$. We denote the trace of a square matrix $A$ by $\operatorname*{tr}(A)$. Let $\mathcal{O}$ be Landau’s symbol; i.e., for a sequence $(x_{k})_{k=0}^{\infty}$, we write $x_{k}=\mathcal{O}(g(k))$ as $k\to\infty$ if there exist a constant $C>0$ and an index $k_{0}$ such that $|x_{k}|\leq C|g(k)|$ for all $k\geq k_{0}$.
In this paper, we focus on Riemannian manifolds (Sakai, 1996) embedded in Euclidean space. Let $M$ be an embedded submanifold of $\mathbb{R}^{d}$. Moreover, let $T_{x}M$ be the tangent space at a point $x\in M$ and $TM$ be the tangent bundle of $M$. Let $0_{x}$ be the zero element of $T_{x}M$. The inner product $\left\langle{\cdot,\cdot}\right\rangle_{2}$ of a Euclidean space $\mathbb{R}^{d}$ induces a Riemannian metric $\left\langle{\cdot,\cdot}\right\rangle_{x}$ of $M$ at $x\in M$ according to $\left\langle{\xi,\eta}\right\rangle_{x}=\left\langle{\xi,\eta}\right\rangle_{2}=\xi^{\top}\eta$ for $\xi,\eta\in T_{x}M\subset T_{x}\mathbb{R}^{d}\cong\mathbb{R}^{d}$. The norm of $\eta\in T_{x}M$ is defined as $\left\lVert{\eta}\right\rVert_{x}=\sqrt{\eta^{\top}\eta}=\left\lVert{\eta}\right\rVert_{2}$. Let $P_{x}:T_{x}\mathbb{R}^{d}\cong\mathbb{R}^{d}\to T_{x}M$ be the orthogonal projection onto $T_{x}M$ (see Absil et al. (2008)). For a smooth map $F:M\to N$ between two manifolds $M$ and $N$, $\mathrm{D}F(x):T_{x}M\to T_{F(x)}N$ denotes the derivative of $F$ at $x\in M$. The Riemannian gradient $\operatorname*{grad}f(x)$ of a smooth function $f:M\to\mathbb{R}$ at $x\in M$ is defined as a unique tangent vector at $x$ satisfying $\left\langle{\operatorname*{grad}f(x),\eta}\right\rangle_{x}=\mathrm{D}f(x)[\eta]$ for any $\eta\in T_{x}M$.
@@ 2.1 Examples
The unit sphere $\mathbb{S}^{d-1}:=\{x\in\mathbb{R}^{d}\mid\left\lVert{x}\right\rVert_{2}=1\}$ is an embedded manifold of $\mathbb{R}^{d}$. The tangent space $T_{x}\mathbb{S}^{n-1}$ at $x\in\mathbb{S}^{d-1}$ is given by $T_{x}\mathbb{S}^{n-1}=\{\eta\in\mathbb{R}^{d}\mid\eta^{\top}x=0\}$. The induced Riemannian metric on $\mathbb{S}^{d-1}$ is given by $\left\langle{\xi,\eta}\right\rangle_{x}=\left\langle{\xi,\eta}\right\rangle_{2}:=\xi^{\top}\eta$ for $\xi,\eta\in T_{x}\mathbb{S}^{d-1}$. The orthogonal projection $P_{x}:\mathbb{R}^{d}\to T_{x}\mathbb{S}^{d-1}$ onto the tangent space $T_{x}\mathbb{S}^{n-1}$ is given by $P_{x}(\eta)=(I_{d}-xx^{\top})\eta$ for $x\in\mathbb{S}^{d-1}$ and $\eta\in T_{x}\mathbb{S}^{d-1}$.
@@
The induced Riemannian metric on $\operatorname*{St}(p,n)$ is given by $\left\langle{\xi,\eta}\right\rangle_{X}=\operatorname*{tr}(\xi^{\top}\eta)$ for $\xi,\eta\in T_{X}\operatorname*{St}(p,n)$. The orthogonal projection onto the tangent space $T_{X}\operatorname*{St}(p,n)$ is given by $P_{X}(\eta)=\eta-X\operatorname*{sym}(X^{\top}\eta)$ for $X\in\operatorname*{St}(p,n)$, $\eta\in T_{X}\operatorname*{St}(p,n)$, where $\operatorname*{sym}(A):=(A+A^{\top})/2$. The Stiefel manifold $\operatorname*{St}(p,n)$ reduces to the orthogonal groups when $n=p$, i.e., $\operatorname*{St}(p,p)=\mathcal{O}_{p}$.
@@
for $[X]\in\operatorname*{Gr}(p,n)$ and $\eta\in T_{[X]}\operatorname*{Gr}(p,n)$. Note that while the Grassmann manifold itself is not an embedded submanifold, in practical implementations, computations involving it is often performed via the Stiefel manifold. This allows the proposed method to be applied. Such an approach was used in previous studies, including Kasai et al. (2019); Hu et al. (2024).